29. 256=16c
31. 400=25n
33. 1/2x=-40
35. 0=3/5t

Answer by 303795(602)

256=16c
c=256/16
=16

400=25n
n=400/25
=16

1/2x=-40
x=-40*2
=-80

0=3/5t
t=0